OpenStudy (jmcs1985):

Find the derivative x^2 + 2tanx

OpenStudy (anonymous):

alright, lets break this into parts can you derivate x^2 and derivate 2tanx? if not I can help

OpenStudy (jmcs1985):

Yup....\[\frac{ d }{dx }x^2 = 2x, \frac{d}{dx}2tanx = 2\sec ^{2}x\]

OpenStudy (jmcs1985):

after chain rule

OpenStudy (anonymous):

right, so now unseparate them. you are basically done

OpenStudy (jmcs1985):

So, \[2x+2\sec^{2}x\]?

OpenStudy (anonymous):

bingo

OpenStudy (jmcs1985):

Awesome, Thanks!